OpenStudy (anonymous):

Simplify the expression the root of negative sixteen all over the quantity of three minus three i plus the quantity of one minus two i. .

OpenStudy (anonymous):

@Hero

Nnesha (nnesha):

\[\huge\rm \frac{ \sqrt{-16} }{ (3-3i)+(1-2i) }\] like this eh?

Nnesha (nnesha):

combine like terms at the denominator factor out the square root of -1 from sqrt{-16} and remember \[\large\rm i = \sqrt{-1}\]

Nnesha (nnesha):

wll continue later...